Please help me! These are two separate problems.
Let f(x)=x+4 and g(x)=x^2 -4.
Find and simplify the expression. (f+g)(2)
(f+g)(2=_________ (Simplify your answer.)
Let f(x)=x^2 -16 and g(x)=4-x.
Find (fg)(-3).
(fg)(-3)=____________

Answer by ewatrrr(24785)


Hi
f(x)=x+4 and g(x)=x^2 -4
(f+g)(x) = x+4 + x^2 - 4 = x^2 + x
(f+g)(x) = x^2 + x
(f+g)(2) = 2^2 + 2 = 6
f(x)=x^2 -16 and g(x)=4-x
(f+g)(x) = x^2 - 16 + 4- x
(f+g)(x) = x^2 -x -12
(f+g)(-3) = (-3)^2 -(-3) -12 = 0